OpenStudy (ragingsquirrel):

****HELP***HELP***HELP*** A rectangular prism is 4 m long, 5 m wide, and has a height of 6 m. What is its surface area? 76 m2 120 m2 148 m2 240 m2

OpenStudy (ragingsquirrel):

Of couse that worked....

OpenStudy (anonymous):

Lateral surface area = \( 2 \times (l+b) \times h \)

OpenStudy (anonymous):

surface area = 2×(l+b)×h

OpenStudy (ragingsquirrel):

What's b?

OpenStudy (anonymous):

Total surface area = Lateral surface area \( + 2 \times l\times b \) l = length b= breadth h = height

OpenStudy (ragingsquirrel):

108 isn't an option...

OpenStudy (ragingsquirrel):

Well, i'm just gunna guess. Thanks for trying

OpenStudy (anonymous):

The answer is \( \boxed{c}\)

OpenStudy (anonymous):

240 m2 did wrong formula ...srry